OpenStudy (anonymous):

find the exponential equation of the line that best describes the provided information. (3,-243) and (5,-19683)

OpenStudy (perl):

technically it is not a line but a curve

OpenStudy (anonymous):

WELCOME TO OPENSTUDY and what do u think it is?

OpenStudy (anonymous):

im not really sure what it is or how to do it honestly.

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

2 points can always be a line

OpenStudy (solomonzelman):

I mean it would be a line with a very small negative slope (using "small" knowing that -4>-10)

OpenStudy (perl):

i think he wants an exponential curve, though

OpenStudy (solomonzelman):

yes, most likely. It is not necesarily (only) a curve though.

OpenStudy (solomonzelman):

\(\large\color{royalblue}{ \displaystyle y=a(b)^x }\) plug in your points, and make the system of equations to solve.

OpenStudy (solomonzelman):

Like: \(\large\color{royalblue}{ \displaystyle -243=a(b)^3 }\) \(\large\color{royalblue}{ \displaystyle -19683=a(b)^5 }\)

OpenStudy (solomonzelman):

then solve the sys of equations for a and b. As you find a and b, re-write the \(\large\color{royalblue}{ \displaystyle y=a(b)^x }\), BUT with the values of a and b that you find.

OpenStudy (solomonzelman):

fro there it should be clear enough....

OpenStudy (anonymous):

okay thank you

OpenStudy (solomonzelman):

sure

OpenStudy (anonymous):

do you mind if i message you?

OpenStudy (solomonzelman):

well, it depends on the content of the message. I do mind "oh please give me the answer, and I will fan you"

OpenStudy (solomonzelman):

But if you ask a question about the problem I answered to, then it would be fair that you ask it in the post, and that I come back to it.

OpenStudy (anonymous):



OpenStudy (solomonzelman):

yes, once you found b, plug it into (i would recommend) the -243=b^3 to find a.

OpenStudy (solomonzelman):

I mean into -243=a(b)^3

OpenStudy (anonymous):

okay.

OpenStudy (solomonzelman):

did you find a?